# A quantum speedup algorithm for TSP based on quantum dynamic programming with very few qubits

### v2

@@ Abstract
The Traveling Salesman Problem (TSP) is a classical NP-hard problem that plays a crucial role in combinatorial optimization. In this paper, we are interested in the quantum search framework for the TSP because it has robust theoretical guarantees. However, we need to first search for all Hamiltonian cycles from a very large solution space, which greatly weakens the advantage of quantum search algorithms. To address this issue, one can first prepare a superposition state of all feasible solutions, and then amplify the amplitude of the optimal solution from it. We propose a quantum algorithm to generate the uniform superposition state of all N-length Hamiltonian cycles as an initial state within polynomial gate complexity based on pure quantum dynamic programming with very few ancillary qubits, which achieves exponential acceleration compared to the previous initial state preparation algorithm. As a result, we realized the theoretical minimum query complexity of quantum search algorithms for a general TSP. Compared to some algorithms that theoretically have lower query complexities but lack practical implementation solutions, our algorithm has feasible circuit implementation.
@@ I Introduction
In this study, we are interested in the quantum search framework for the TSP because it has robust theoretical guarantees. Two realizable algorithms in this direction can serve as references. Zhu et al. designed and improved the Hamiltonian cycle detection (HCD) oracle to determine valid cycles using quantum circuits. Combining the HCD oracle and the quantum phase estimation, they proposed a quantum algorithm for TSP based on GAS with a query complexity of $O(\sqrt{2^{\lceil\log d\rceil N}})$, where $d$ is the maximum degree of the graph [29]. Some researchers [30] realized the importance of preparing the initial state or the superposition state encoding the feasible solutions. They proposed a two-step quantum search (TSQS) algorithm based on higher-order unconstrained binary optimization (HOBO) representation. The first step amplifies the amplitude of feasible solutions to prepare a uniform superposition state, from which the second step determines the optimal solution. The query complexity of the entire algorithm is $O(\sqrt{N!})$, which is close to the minimum query complexity of quantum search algorithms for a general TSP.
@@ IV Discussion
We have transformed the TSP into a minimum value search problem, in which there are two problems. One is how to provide a good Grover’s iteration count setting scheme when the number of the optimal solutions is unknown, which can be solved by referring to [34]. The other is how to search for the minimum value by executing our algorithm multiple times to update the threshold $C_{T}$, which can be solved by referring to [33, 34]. After these improvements, the final algorithm still maintains the query complexity of $O(\sqrt{(N-1)!})$ (simply adding a constant factor). The complete algorithm steps, which are the general steps of Grover’s algorithm for finding the minimum value (just change the number of feasible solutions), are as follows:
@@
Query complexity. Owing to the varying complexity using different decomposition methods of gates and encoding methods of data, when comparing the complexity of algorithms we generally use query complexity, that is, the number of execution times of oracle marking the target states. In Table 3, we list the comparison results with excellent realizable quantum search algorithms for the TSP.
[b] algorithm query complexity GAS(HCD) [29] $O(\sqrt{2^{\lceil\log d\rceil N}})$ TSQS(HOBO) [30] $O(\sqrt{N!})$ Our algorithm $O(\sqrt{(N-1)!})$
Table 3: Comparison Results on query complexity. $d$ is the maximum degree of the graph.

### v4

@@ Abstract
The Traveling Salesman Problem (TSP) is a classical NP-hard problem that plays a crucial role in combinatorial optimization. In this paper, we are interested in the quantum search framework for the TSP because it has robust theoretical guarantees. However, we need to first search for all Hamiltonian cycles from a very large solution space, which greatly weakens the advantage of quantum search algorithms. To address this issue, one can first prepare a superposition state of all feasible solutions, and then amplify the amplitude of the optimal solution from it. We propose a quantum algorithm to generate the uniform superposition state of all $N$-length Hamiltonian cycles as an initial state within polynomial gate complexity based on pure quantum dynamic programming with very few ancillary qubits, which achieves exponential acceleration compared to the previous initial state preparation algorithm. As a result, we realized the theoretical minimum query complexity of quantum search algorithms for a general TSP. Compared to some algorithms that theoretically have lower query complexities but lack practical implementation solutions, our algorithm has feasible circuit implementation. Our work provides a meaningful research case on how to fully utilize the structures of specific problems to unleash the acceleration capability of the quantum search algorithms.
@@ I Introduction
In this study, we are interested in the quantum search framework for the TSP because it has robust theoretical guarantees. Two realizable algorithms in this direction can serve as references. Zhu et al. designed and improved the Hamiltonian cycle detection (HCD) oracle to determine valid cycles using quantum circuits. Combining the HCD oracle and the quantum phase estimation, they proposed a quantum algorithm for TSP based on GAS with a query complexity of $O(\sqrt{2^{\lceil\log d\rceil N}})$, where $d$ is the maximum degree of the graph [29]. And this complexity is approximately equal to $O(\sqrt{N^{N}})$ for $d\sim O(N)$ under the statistical significance of random instances. Some researchers [30] realized the importance of preparing the initial state or the superposition state encoding the feasible solutions. They proposed a two-step quantum search (TSQS) algorithm based on higher-order unconstrained binary optimization (HOBO) representation. The first step amplifies the amplitude of feasible solutions to prepare a uniform superposition state, from which the second step determines the optimal solution. The query complexity of the entire algorithm is $O(\sqrt{N!})$, which is close to the minimum query complexity of quantum search algorithms for a general TSP. However, they need the query complexity of $O(\sqrt{2^{N\log N}/N!})\approx O(e^{N/2}/N^{1/4})$ (according to Sterling’s approximation) to prepare the initial state.
@@ IV Discussion
We have transformed the TSP into a minimum value search problem, in which there are two problems. One is how to provide a good Grover’s iteration count setting scheme when the number of the optimal solutions is unknown. The other is how to search for the minimum value by executing our algorithm multiple times to update the threshold $C_{T}$. These two problems can be solved by applying quantum exponential searching [33, 34]. After these improvements, the final algorithm still maintains the query complexity of $O(\sqrt{(N-1)!})$ (simply adding a constant factor). The complete algorithm steps, which are the general steps of Grover’s algorithm for finding the minimum value (just change the number of feasible solutions), are as follows:
@@
Query complexity. Owing to the varying complexity using different decomposition methods of gates and encoding methods of data, when comparing the complexity of algorithms we generally use query complexity, that is, the number of execution times of oracle marking the target states. In Table 3, we list the comparison results with excellent realizable quantum search algorithms for the TSP. We compare our algorithm with GAS(HCD) in the average sense. And $O(\sqrt{(N-1)!})$ is the theoretical minimum query complexity of quantum search algorithms for a general TSP because we have to find the shortest one among $(N-1)!$ HCs when we consider random unknown input instances. One may note that for sparse instances with very small $d$, GAS(HCD) has lower query complexity, while what we care about are instances on the general complete graphs, which contain situations of sparse graphs by setting weights on certain edges to 0.
[b]
Table 3: Comparison Results on query complexity. $d$ is the maximum degree of the graph and $d\sim O(N)$ under the statistical significance of random instances.
algorithm query complexity
GAS(HCD) [29] $O(\sqrt{2^{\lceil\log d\rceil N}})\approx O(\sqrt{N^{N}})$
TSQS(HOBO) [30] $O(\sqrt{N!})$
Our algorithm $O(\sqrt{(N-1)!})\approx O(\sqrt{N^{N}/e^{N}})$